OpenStudy (ashy98):

1 subtracted from x is 3 What is x ???

OpenStudy (anonymous):

4

OpenStudy (shamil98):

1 subtracted from x = x-1 is means equal so x-1= 3 + 1 + 1 x =4

OpenStudy (ashy98):

thanks (: